# Primordial Black Holes as Seeds for Extremely Overmassive AGN Observed by JWST

### v1

@@ Abstract
facilities: Lonestar6 and Stampede3 (TACC)software: astropy (2013A&A...558A..33A; 2018AJ....156..123A; 2022ApJ...935..167A), Colossus (Diemer2018ApJCOLOSSUS)
@@ 1 Introduction
Since the launch of the James Webb Space Telescope (JWST), an unexpected population of massive supermassive black holes (SMBHs) has been revealed at very high redshifts ($z\gtrsim 7$(e.g., Goulding2023ApJ...955L..24G; Larson_2023_BH; Bogdan:2023UHZ1; Greene2024; Kovacs2024ApJ...965L..21K; Maiolino2024A&A; Natarajan:2023UHZ1; GHZ9Napolitano2025ApJ...989...75N). Among these early systems, a class of compact, extremely red, and spectroscopically peculiar sources, commonly referred to as “little red dots” (LRDs), has emerged as a distinct population (e.g., Labbe2023Natur.616..266L; KokorvLRD2024ApJ...968...38K; Leung2024:LRDarXiv; KocevskiLRD2025ApJ...986..126K; Taylor2025ApJ...986..165T; LabbeLRD2025ApJ...978...92L). These newly-discovered objects are found to host central black holes (BHs) that are often ‘overmassive’ relative to their stellar components, with black-hole-to-stellar mass ratios exceeding the local relation (e.g. PacucciLRD2023ApJ...957L...3P; Inayoshi2024ApJ...966..164I). Among those discoveries, one of the most striking examples is the JWST source Abell 2744–QSO1, whose inferred BH mass ($M_{\rm BH}\sim 5\times 10^{7}\,M_{\odot}$), lack of significant stellar contributions in its kinematics and spectra ($M_{\star}\lesssim 2\times 10^{7}\ \rm M_{\odot}$), and extremely low metallicity ($Z\lesssim 0.01\,Z_{\odot}$) at $z\sim 7$ pose severe challenges to conventional models of SMBH formation and coevolution with its host galaxy (FurtakAbell27442023ApJ; BlackTHUNDER2025arXiv; Maiolino2025arXiv250522567M; QSO1Direct2025arXiv250821748J).
@@
Traditional light-seed scenarios, typically related to stellar remnants of Population III (Pop III) stars (see reviews in, e.g., Smith2019:BHreview; Inayoshi:2020), struggle to reproduce both the high inferred accretion rates and the black-hole-to-stellar mass ratios implied by JWST observations (e.g., Milosavljevic2009; Jeon2023; Bogdan:2023UHZ1). These models generally require sustained super-Eddington accretion over extended periods (e.g., Alexander2014Sci...345.1330A; Jeon2025ApJ...979..127J). An alternative pathway is the formation of massive BHs through the direct collapse of pristine, metal-poor gas with minimal fragmentation (Loeb1994ApJ...432...52L; BrommDCBH2003ApJ...596...34B; Begelman2006:DCBH; Lodato:2006DCBH). Such conditions typically demand a nearby source of intense soft-UV, Lyman–Werner (LW) radiation to suppress molecular hydrogen cooling (e.g., AgarwalDCBH2014MNRAS.443..648A; Habouzit2016; OBrennan2025), which is absent in the vicinity of Abell 2744–QSO1 (QSO1Direct2025arXiv250821748J).
@@
Another explanation involves the presence of primordial black holes (PBHs) that predate galaxy formation (for analytical work see, e.g., DuchtingPBH2004PhRvD..70f4015D; Bernal:2017nec; DeLuca2023PhRvL.130q1401D; Dayal2024A&A...690A.182D). PBHs, hypothesized to form from the collapse of density fluctuations shortly after the Big Bang (Zeldovich1967SvA....10..602Z; hawking1971gravitationally; Carr1975ApJ...201....1C; Belotsky2019; Escriva2022), may constitute only a small fraction of the dark matter, yet can profoundly influence cosmic evolution (for a review, see Carr2020ARNPS..70..355C). On larger scales, depending on their mass spectrum and abundance, PBHs can enhance primordial density fluctuations and seed the formation of dark matter halos (e.g., Mack2007ApJ; Ricotti2007ApJI; Ricotti2008ApJII; Carr2018MNRAS.478.3756C; Inman2019PhRvD.100h3528I; Su2023; Colazo2024; Zhang:2024PBH), thus accelerating early structure formation (Meszaros1975A&A....38....5M; Afshordi2003ApJ...594L..71A; Kashlinsky2021PhRvL.126a1101K; Cappelluti2022ApJ; Boyuan2022ApJ). Their early accretion can heat and ionize the intergalactic medium (IGM), thus modifying cosmic radiation backgrounds (e.g., Ali-Haimoud2017PhRvD; Deluca2020JCAP...06..044D; Lu2021; Takhistov2022; Ziparo2022; Zhang2024MNRAS.528..180Z; Casanueva-Villarreal2024), and potentially trigger the first episodes of star formation around PBH-seeded halos (Boyuan2022MNRAS.514.2376L; Boyuan2023arXiv231204085L). Recent studies further suggest that massive PBHs ($\gtrsim 10^{6}\,M_{\odot}$) can induce the formation of bound stellar clusters (Zhang2025PBHStar) or secondary massive BHs (Zhang2025pbhdcbh), giving rise to systems reminiscent of the LRDs discovered by JWST (DayalPBH2025arXiv250608116D; Matteri2025arXiv250318850M). However, most previous simulations have included only BH accretion and feedback while ignoring stellar feedback (or vice versa), leaving open how PBHs coevolve with PBH-seeded galaxies in the early Universe.
@@
In this work, we integrate the legacy stellar feedback model (Jaacks2018; Jaacks2019; Liu2020MNRAS.497.2839L; Liu2020) into our cosmological PBH simulation framework. This approach enables, for the first time, a combined treatment of both BH and stellar feedback in the early growth of PBH-seeded systems. Our simulations follow the coupled evolution of dark matter, baryons and stellar components from $z\simeq 1100$ to $z\simeq 7$, reproducing the physical conditions of Abell 2744–QSO1 and potentially other JWST LRDs, aiming to assess whether heavy PBH seeds can naturally account for the formation of these extreme high-redshift systems.
@@
In Section 2, we describe the numerical setup for the simulation, focusing on the recipes for star formation and stellar feedback. From the simulated results, the detailed galaxy assembly and metal enrichment history around the central PBH is discussed in Section 3. The limitations and caveats are addressed in Section 4, and we present our conclusions in Section 5.
@@
Table 1: Summary of key parameters and main results. $L$ is the size of the box in comoving units, $z_{\rm ini}$ the initial redshift where the simulation starts, and $N_{\rm eff}$ the total number of particles within the simulation box. $\epsilon_{r}$ is the thermal coupling efficiency of BH feedback, FULL_SF a flag to control whether we consider the full stellar feedback recipe or not, whereas $\epsilon_{\rm DM}$, $\epsilon_{\rm gas}$, and $\epsilon_{\star}$ represent the (co-moving) softening lengths for PDM, gas, and stellar particles, respectively. Finally, $z_{\rm col}$ is the redshift when star formation begins around the central PBH, and $M_{\star}$ denotes the total mass of stars formed in the PBH-hosting halo, evaluated by the end of the simulation ($z\sim 7$).
@@
Run $L$ [Mpc/h] $z_{\rm ini}$ $N_{\rm eff}$ $\epsilon_{\rm r}$ FULL_SF $\epsilon_{\rm DM}[\rm kpc/h]$ $\epsilon_{\rm gas}[\rm kpc/h]$ $\epsilon_{\star}[\rm kpc/h]$ $z_{\rm col}$ $M_{\star}[\mathrm{M}_{\odot}]$
PBH_DMonly 1 3400 $256^{3}$ - 0.05 - - - - -
PBH_M5e7_fd005 1 1100 $2\times 256^{3}$ 0.005 0.05 0.05 0.005 9.64 $2.0\times 10^{7}$
PBH_SF_M5e7_fd005 1 1100 $2\times 256^{3}$ 0.005 0.05 0.05 0.005 8.20 $7.7\times 10^{5}$
2 Methodology
@@
The simulation is performed using the GIZMO code (Hopkins2015MNRAS.450...53H), which couples the Lagrangian meshless finite-mass (MFM) hydrodynamics solver (with $N_{\rm ngb}=32$ neighbors) to the parallel Tree+PM gravity solver of GADGET-3 (springel2005cosmological). The non-equilibrium primordial chemistry and cooling network includes 12 species, following Bromm2002ApJ...564...23B and Johnson2006MNRAS.366..247J, supplemented by C II, O I, Si II and Fe II line cooling (Jaacks2018), with initial abundances at $z\simeq 1100$ set by Galli2013ARA&A..51..163G. The Planck18 cosmological parameters were adopted throughout (Plank2020A&A...641A...6P): $\Omega_{\rm m}=0.3111$, $\Omega_{\rm b}=0.04897$, $h=0.6776$, $\sigma_{8}=0.8102$, $n_{\rm s}=0.9665$.
@@
We describe the initial conditions and general numerical setup in Section 2.1, and the stellar formation and feedback prescriptions in Section 2.2. The relevant parameters for all runs are summarized in Table 1.
2.1 Initial Conditions
@@
To reproduce the observed properties of Abell 2744–QSO1, we focus on the cosmic evolution surrounding an isolated PBH. The simulation box has a comoving side length of $L=1\,h^{-1}\,\mathrm{Mpc}$ and contains $N_{\rm PDM}=256^{3}$ dark matter particles. The simulation begins at $z\simeq 3400$, where we place a BH of mass $5\times 10^{7}\,M_{\odot}$ at the center of the box [^1] In this work, we simulate structure formation around an isolated PBH, motivated by the inferred mass of Abell 2744–QSO1 (QSO1Direct2025arXiv250821748J). The chosen box size limits the effective PBH mass fraction to $f_{\rm PBH}\lesssim 8\times 10^{-4}$, consistent with the mass function assumed by Carr2021PDU....3100755C. However, such a massive PBH could also originate from mergers of initially clustered lower-mass PBHs (Zhang:2024PBH; ZhangPBHCluster2025arXiv250707171Z)., generated using the MUSIC code (hahn2011multi). The numerical procedure for embedding the PBH and perturbing the surrounding particle dark matter (PDM) field is implemented using the PHANTOM code (zhang_2025_17025634), following the formalism in Ali-Haimoud2017PhRvD; Inman2019PhRvD.100h3528I; Boyuan2022MNRAS.514.2376L; Zhang:2024PBH, providing the initial condition for the PBH_DMonly run. We first evolve this run from $z\simeq 3400$ to the recombination epoch ($z\simeq 1100$) to establish the PDM distribution around the central PBH. We then add a uniform baryonic component to initialize the full hydrodynamical runs, resulting in a total of $N=2\times 256^{3}$ particles for both the PBH_SF_M5e7_fd005 and PBH_M5e7_fd005 runs with and without stellar feedback, respectively. The resulting gas and dark matter particle masses are $1.2\times 10^{3}\,M_{\odot}$ and $5.1\times 10^{3}\,M_{\odot}$.
@@
BH accretion and feedback are modeled using the same subgrid prescription as in Zhang2025PBHStar which self-consistently tracks the growth of the central PBH through Bondi-Hoyle-Lyttleton accretion with radiative/thermal feedback. The feedback is implemented by depositing energy into the surrounding gas according to a fiducial thermal radiation coupling efficiency of $\epsilon_{r}=\delta E_{\bullet,\rm inject}/(L_{\bullet,\rm acc}\delta t)=0.005$  [^2] The adopted feedback coefficient is also chosen to reproduce the observed properties of the QSO-1 system, allowing for moderate star formation consistent with its inferred stellar content. We refer the reader to our previous work (Zhang2025PBHStar) and to a forthcoming study for a more comprehensive discussion of accretion feedback and the assumed seed mass., where $\delta E_{\bullet,\rm inject}$ represents the thermal energy injected into the ambient medium, and $L_{\bullet,\rm acc}$ is the accretion luminosity. Both runs are evolved from $z\simeq 1100$ to $z\simeq 7$, corresponding to the observed redshift of Abell 2744–QSO1 (BlackTHUNDER2025arXiv).
2.2 Star Formation and Stellar Feedback

### v2

@@ Abstract
facilities: Lonestar6 and Stampede3 (TACC)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Colossus (Diemer, 2018)
@@ I Introduction
Since the launch of the James Webb Space Telescope (JWST), an unexpected population of massive supermassive black holes (SMBHs) has been revealed at very high redshifts ($z\gtrsim 7$(e.g., Goulding et al., 2023; Larson et al., 2023; Bogdán et al., 2024; Greene et al., 2024; Kovács et al., 2024; Maiolino et al., 2024; Natarajan et al., 2024; Napolitano et al., 2025). Among these early systems, a class of compact, extremely red, and spectroscopically peculiar sources, commonly referred to as “little red dots” (LRDs), has emerged as a distinct population (e.g., Labbé et al., 2023; Kokorev et al., 2024; Leung et al., 2025; Kocevski et al., 2025; Taylor et al., 2025; Labbe et al., 2025). These newly-discovered objects are found to host central black holes (BHs) that are often ‘overmassive’ relative to their stellar components, with black-hole-to-stellar mass ratios exceeding the local relation (e.g. Pacucci et al., 2023; Inayoshi et al., 2024). Among those discoveries, one of the most striking examples is the JWST source Abell 2744–QSO1, whose inferred BH mass ($M_{\rm BH}\sim 5\times 10^{7}\,M_{\odot}$), lack of significant stellar contributions in its kinematics and spectra ($M_{\star}\lesssim 2\times 10^{7}\ \rm M_{\odot}$), combined with extremely low metallicity ($Z\lesssim 0.01\,Z_{\odot}$) and low accretion efficiency ($f_{\rm Edd}\sim 0.01-0.03$) at $z\sim 7$ pose severe challenges to conventional models of SMBH formation and coevolution with its host galaxy (Furtak et al., 2023; Ji et al., 2025; Maiolino et al., 2025; Juodžbalis et al., 2025).
@@
Traditional light-seed scenarios, typically related to stellar remnants of Population III (Pop III) stars (see reviews in, e.g., Smith and Bromm, 2019; Inayoshi et al., 2020), struggle to reproduce both the high inferred accretion rates and the black-hole-to-stellar mass ratios implied by JWST observations (e.g., Milosavljević et al., 2009; Jeon et al., 2023; Bogdán et al., 2024). These models generally require sustained super-Eddington accretion over extended periods (e.g., Alexander and Natarajan, 2014; Jeon et al., 2025a). An alternative pathway is the formation of massive BHs through the direct collapse of pristine, metal-poor gas with minimal fragmentation (Loeb and Rasio, 1994; Bromm and Loeb, 2003; Begelman et al., 2006; Lodato and Natarajan, 2006). Such conditions typically demand a nearby source of intense soft-UV, Lyman–Werner (LW) radiation to suppress molecular hydrogen cooling (e.g., Agarwal et al., 2014; Habouzit et al., 2016; O’Brennan et al., 2025), which is absent in the vicinity of Abell 2744–QSO1 (Juodžbalis et al., 2025).
@@
Another explanation involves the presence of primordial black holes (PBHs) that predate galaxy formation (for analytical work see, e.g., Düchting, 2004; Bernal et al., 2018; De Luca et al., 2023; Dayal, 2024; De Luca et al., 2025)  [^1] A few alternative formation scenarios have been proposed, such as core collapse of strongly self-interacting dark matter (e.g., Jiang et al., 2026) or direct collapse of gas at $z\gtrsim 100$ from the enhanced power spectrum at small scales (e.g., Qin et al., 2025). PBHs, hypothesized to form from the collapse of density fluctuations shortly after the Big Bang (Zel’dovich and Novikov, 1967; Hawking, 1971; Carr, 1975; Belotsky et al., 2019; Escrivà, 2022), may constitute only a small fraction of the dark matter, yet can profoundly influence cosmic evolution (for a review, see Carr and Kühnel, 2020). On larger scales, depending on their mass spectrum and abundance, PBHs can enhance primordial density fluctuations and seed the formation of dark matter halos (e.g., Mack et al., 2007; Ricotti, 2007; Ricotti et al., 2008; Carr and Silk, 2018; Inman and Ali-Haïmoud, 2019; Su et al., 2023; Colazo et al., 2024; Zhang et al., 2024a), thus accelerating early structure formation (Meszaros, 1975; Afshordi et al., 2003; Kashlinsky, 2021; Cappelluti et al., 2022; Liu and Bromm, 2022). Their early accretion can heat and ionize the intergalactic medium (IGM), thus modifying cosmic radiation backgrounds (e.g., Ali-Haïmoud and Kamionkowski, 2017; De Luca et al., 2020; Lu et al., 2021; Takhistov et al., 2022; Ziparo et al., 2022; Zhang et al., 2024b; Casanueva-Villarreal et al., 2024), and potentially trigger the first episodes of star formation around PBH-seeded halos (Liu et al., 2022; Liu and Bromm, 2023). Recent studies further suggest that massive PBHs ($\gtrsim 10^{6}\,M_{\odot}$) can induce the formation of bound stellar clusters (Zhang et al., 2025b) or secondary massive BHs (Zhang et al., 2025c), giving rise to systems reminiscent of the LRDs discovered by JWST (Dayal and Maiolino, 2025; Matteri et al., 2025). However, most previous simulations have included only BH accretion and feedback while ignoring stellar feedback (or vice versa), leaving open how PBHs coevolve with PBH-seeded galaxies in the early Universe.
@@
In this work, we integrate the legacy stellar feedback model (Jaacks et al., 2018, 2019; Liu and Bromm, 2020b, a) into our cosmological PBH simulation framework. This approach enables, for the first time, a combined treatment of both BH and stellar feedback in the early growth of PBH-seeded systems. Our simulations follow the coupled evolution of dark matter, baryons and stellar components from $z\simeq 1100$ to $z\simeq 7$, reproducing the physical conditions of Abell 2744–QSO1 and potentially other JWST LRDs, aiming to assess whether heavy PBH seeds can naturally account for the formation of these extreme high-redshift systems.
@@
In Section II, we describe the numerical setup for the simulation, focusing on the recipes for star formation and stellar feedback. From the simulated results, the detailed galaxy assembly and metal enrichment history around the central PBH is discussed in Section III. The limitations and caveats are addressed in Section IV, and we present our conclusions in Section V.
@@
Table 1: Summary of key parameters and main results. $z_{\rm ini}$ the initial redshift where the simulation starts, and $N_{\rm eff}$ the total number of particles within the simulation box of size $L=1\,h^{-1}\,\mathrm{Mpc}$. $\epsilon_{r}$ is the thermal coupling efficiency of BH feedback, $\langle f_{\rm edd}\rangle\equiv\langle\dot{m}_{\rm BH}/\dot{m}_{\rm edd}\rangle$ is the time-averaged BH accretion rate in units of the Eddington limit, FULL_SF a flag to control whether we consider the full stellar feedback recipe or not, whereas $\epsilon_{\rm DM}$, $\epsilon_{\rm gas}$, and $\epsilon_{\star}$ represent the (co-moving) softening lengths for PDM, gas, and stellar particles, respectively. Finally, $z_{\rm col}$ is the redshift when star formation begins around the central PBH, and $M_{\star}$ denotes the total mass of stars formed in the PBH-hosting halo, evaluated by the end of the simulation ($z\sim 7$).
@@
Run $z_{\rm ini}$ $N_{\rm eff}$ $\epsilon_{\rm r}$ $\langle f_{\rm edd}\rangle$ FULL_SF $\epsilon_{\rm DM}\small[\rm kpc/h]$ $\epsilon_{\rm gas}\small[\rm kpc/h]$ $\epsilon_{\star}\small[\rm kpc/h]$ $z_{\rm col}$ $M_{\star}\small[\mathrm{M}_{\odot}]$
PBH_DMonly 3400 $256^{3}$ - - - 0.05 - - - -
PBH_M5e7_fd005 1100 $2\times 256^{3}$ 0.005 0.0076 0.05 0.05 0.005 9.64 $2.0\times 10^{7}$
PBH_SF_M5e7_fd005 1100 $2\times 256^{3}$ 0.005 0.0081 0.05 0.05 0.005 8.20 $7.7\times 10^{5}$
PBH_SF_M5e7_fd0025 1100 $2\times 256^{3}$ 0.0025 0.013 0.05 0.05 0.005 8.58 $3.2\times 10^{5}$
PBH_SF_M5e7_fd01 1100 $2\times 256^{3}$ 0.01 0.0047 0.05 0.05 0.005 12.80 $1.9\times 10^{6}$
II Methodology
@@
The simulation is performed using the GIZMO code (Hopkins, 2015), which couples the Lagrangian meshless finite-mass (MFM) hydrodynamics solver (with $N_{\rm ngb}=32$ neighbors) to the parallel Tree+PM gravity solver of GADGET-3 (Springel, 2005). The non-equilibrium primordial chemistry and cooling network includes 12 species, following Bromm et al. (2002) and Johnson and Bromm (2006), supplemented by C II, O I, Si II and Fe II line cooling (Jaacks et al., 2018), with initial abundances at $z\simeq 1100$ set by Galli and Palla (2013). The Planck18 cosmological parameters were adopted throughout (Planck Collaboration et al., 2020): $\Omega_{\rm m}=0.3111$, $\Omega_{\rm b}=0.04897$, $h=0.6776$, $\sigma_{8}=0.8102$, $n_{\rm s}=0.9665$.
@@
We describe the initial conditions and general numerical setup in Section II.1, and the stellar formation and feedback prescriptions in Section II.2. The relevant parameters for all runs are summarized in Table 1.
II.1 Initial Conditions
@@
To reproduce the observed properties of Abell 2744–QSO1, we focus on the cosmic evolution surrounding an isolated PBH. The simulation box has a comoving side length of $L=1\,h^{-1}\,\mathrm{Mpc}$ and contains $N_{\rm PDM}=256^{3}$ dark matter particles. The simulation begins at $z\simeq 3400$, where we place a BH of mass $5\times 10^{7}\,M_{\odot}$ at the center of the box  [^2] In this work, we simulate structure formation around an isolated PBH, motivated by the inferred mass of Abell 2744–QSO1 (Juodžbalis et al., 2025). The chosen box size limits the effective PBH mass fraction to $f_{\rm PBH}\lesssim 8\times 10^{-4}$, which is broadly compatible with formation scenarios based on early-Universe phase transitions (Carr et al., 2021a). However, such a massive PBH could also originate from mergers of initially clustered lower-mass PBHs (Zhang et al., 2024a, 2025a)., generated using the MUSIC code (Hahn and Abel, 2011). The numerical procedure for embedding the PBH and perturbing the surrounding particle dark matter (PDM) field is implemented using the PHANTOM code (Zhang et al., 2025d), following the formalism in Ali-Haïmoud and Kamionkowski (2017); Inman and Ali-Haïmoud (2019); Liu et al. (2022); Zhang et al. (2024a), providing the initial condition for the PBH_DMonly run. We first evolve this run from $z\simeq 3400$ to the recombination epoch ($z\simeq 1100$) to establish the PDM distribution around the central PBH  [^3] The PBH is treated as a freely moving collisionless particle providing the central gravitational potential; numerically, it is evolved like any other particle except for a boosted drag term during accretion to minimize BH wandering. To be specific, we apply a drag force $\vec{F}=-\dot{m}_{\rm edd}\vec{v}_{\rm rel}$ to the BH, given its Eddington accretion rate $\dot{m}_{\rm edd}$ and velocity $\vec{v}_{\rm rel}$ relative to gas. Consequently, the BH-gas relative velocity remains well below the local sound speed throughout the simulation, rendering its contribution to the accretion rate subdominant.. We then add a uniform baryonic component to initialize the full hydrodynamical runs, resulting in a total of $N=2\times 256^{3}$ particles for both the PBH_SF_M5e7_fd005 and PBH_M5e7_fd005 runs with and without stellar feedback, respectively. The resulting gas and dark matter particle masses are $1.2\times 10^{3}\,M_{\odot}$ and $5.1\times 10^{3}\,M_{\odot}$.
@@
BH accretion and feedback are modeled using the same subgrid prescription as in Zhang et al. (2025b) which self-consistently tracks the growth of the central PBH through Bondi-Hoyle-Lyttleton accretion with radiative/thermal feedback. The feedback is implemented by depositing energy into the surrounding gas according to a fiducial thermal radiation coupling efficiency of $\epsilon_{r}=\delta E_{\bullet,\rm inject}/(L_{\bullet,\rm acc}\delta t)=0.005$  [^4] The adopted feedback coefficient is also chosen to reproduce the observed properties of the QSO-1 system, allowing for moderate star formation consistent with its inferred stellar content. We have also included two additional runs ($\epsilon_{r}=0.0025$, PBH_SF_M5e7_fd0025 and $\epsilon_{r}=0.01$, PBH_SF_M5e7_fd01) by varying this parameter (see Table 1). In this work, we do not explicitly model UV/LW radiation from BH accretion, given the limited understanding of the BH intrinsic SED in the early universe (see, e.g., Greene et al., 2026). We refer the reader to our previous work (Zhang et al., 2025b, c), and to a forthcoming study for a more comprehensive discussion of accretion feedback and the assumed seed mass., where $\delta E_{\bullet,\rm inject}$ represents the thermal energy injected into the ambient medium, and $L_{\bullet,\rm acc}$ is the accretion luminosity. Both runs are evolved from $z\simeq 1100$ to $z\simeq 7$, corresponding to the observed redshift of Abell 2744–QSO1 (Ji et al., 2025).
II.2 Star Formation and Stellar Feedback